### 10 Jan 2024 v1

# Knowledge Sharing in Manufacturing using Large Language Models: User Evaluation and Model Benchmarking

Abstract.
Managing knowledge efficiently is crucial for organizational success. In manufacturing, operating factories has become increasing knowledge-intensive putting strain on the factory’s capacity to train and support new operators. In this paper, we introduce a Large Language Model (LLM)-based system designed to use the extensive knowledge contained in factory documentation. The system aims to efficiently answer queries from operators and facilitate the sharing of new knowledge. To assess its effectiveness, we conducted an evaluation in a factory setting. The results of this evaluation demonstrated the system’s benefits; namely, in enabling quicker information retrieval and more efficient resolution of issues. However, the study also highlighted a preference for learning from a human expert when such an option is available. Furthermore, we benchmarked several closed and open-sourced LLMs for this system. GPT-4 consistently outperformed its counterparts, with open-source models like StableBeluga2 trailing closely, presenting an attractive option given its data privacy and customization benefits. Overall, this work offers preliminary insights for factories considering using LLM-tools for knowledge management.
1. Introduction
Managing knowledge has become invaluable in today’s manufacturing industry, and technology plays an ever-increasing role. Factories already generated a wealth of knowledge in the form of issue reports and machine documentation; however, up till now, these sources have been challenging to harness using AI. With the advent of Large Language Models, such as GPT-4, the potential to use technology to support operators in knowledge-intensive tasks has leaped forward. However, using LLMs comes with many (unknown) practical and socio-technical risks and challenges. Furthermore, a generic LLM has limited use in a factory, where specific knowledge is needed. As such, an LLM-powered tool for manufacturing would require fine-tuning the model or including context information in its prompt (e.g., sections from a machine manual). We built a fully functional LLM-powered tool for this work that finds relevant sections from uploaded documents and issue reports to answer operator queries. We conducted a user study at a factory to evaluate the usability and potential of the system. Furthermore, we benchmarked several LLMs for the system, including closed and open source options. LLMs are typically benchmarked by the factuality of their response to general knowledge questions, resulting in LLM leaderboards111https://huggingface.co/spaces/HuggingFaceH4/open_llm_leaderboard. However, the ability to answer general knowledge questions is less relevant to a manufacturing environment where very specific, dynamic knowledge is required. Our findings could help decide how factories use this technology in the future.
1.1. Industry 5.0
Industry 5.0, the latest phase of industrial development, places human beings at the forefront of manufacturing processes, emphasizing their skills, creativity, and problem-solving abilities (Alves et al., 2023). This human-centered approach addresses the dehumanization of work, which has been a challenge in previous industrial revolutions. Industry 5.0 seeks to create a more fulfilling and productive work environment by empowering the workforce.
One of the fundamental problems that Industry 5.0 addresses is the alienation and dissatisfaction experienced by workers as a result of automation and machine dominance (Alves et al., 2023). The rapid advancement of technology has often led to a perception that humans are mere operators, overshadowed by the efficiency and precision of machines. However, in Industry 5.0, humans regain their significance in manufacturing. By integrating advanced technologies such as robotics and artificial intelligence, Industry 5.0 enables humans to collaborate with machines, leveraging their complementary strengths. This collaboration enhances productivity and efficiency and acknowledges the value of human contributions.
Human-centered manufacturing in Industry 5.0 focuses on providing a work environment that nurtures individuals’ creativity and problem-solving capabilities (Maddikunta et al., 2022). It encourages workers to think critically, innovate, and continuously learn. With machines handling repetitive and mundane tasks, human workers can dedicate their time and energy to more complex and intellectually stimulating activities. This shift could enhance job satisfaction and promote personal and professional growth, as workers could acquire new skills and engage in higher-level decision-making (Alves et al., 2023), (Xu et al., 2021). Emphasis on human-machine collaboration and the continuous emergence and refinement of technology increase the need for adequate human-computer interaction (Brückner et al., 2023). One of the approaches to address this topic is using conversational AI to assist humans in manufacturing (Wellsandt et al., 2021).
1.2. Large-Language Model-Powered Tools for Knowledge-Intensive Scenarios
The pre-training of LLMs on large-scale, mixed-source corpora allows them to capture extensive knowledge (Zhao et al., 2023), and demonstrate proficiency in interpreting complex information (Jawahar et al., 2019), generating general reasoning and insights (Wei et al., 2022), and aiding knowledge-intensive decision-making in various contexts. Consequently, researchers have been exploring LLMs adoption in both general and domain-specific tasks.
However, despite their advantages, the responses generated by LLMs in the mentioned scenarios may encounter two potential issues: (1) outdated information originating from the model’s pre-training date, and (2) inaccuracies in factual representation, known as “hallucination”(Bang et al., 2023; Zhao et al., 2023). To address these challenges and leverage the capabilities of LLMs in domain-specific knowledge-intensive tasks, chain-of-thought and few-shot prompting (Brown et al., 2020; Gao et al., 2021) techniques are utilized.
This section explores LLM-based tools designed to acquire information and support decision making in both non-manufacturing and manufacturing settings.
Using few-shot prompting to retrieve information across diverse topics, (Semnani et al., 2023) introduced an open-domain LLM-powered chatbot called WikiChat. WikiChat utilizes a 7-stage pipeline of few-shot prompted LLM that suggests facts verified against Wikipedia, retrieves additional up-to-date information, and generates coherent responses. They used a hybrid human-and-LLM method to evaluate the chatbot on different topics for factuality, alignment with real-worth truths and verifiable facts, and conversationality, a compound metric that scores how informational, natural, non-repetitive, and temporally correct the response is. Their solution significantly outperforms GPT-3.5 in factuality, with an average improvement of 24.4% while staying on par in conversationality. Others have explored the capabilities of LLMs in domain-specific tasks such as extracting structured data from unstructured healthcare texts (Tang et al., 2023), providing medical advice (Nov et al., 2023), simplifying radiology reports (Jeblick et al., 2022), Legal Judgement Prediction from multilingual legal documents (Trautmann et al., 2022), and scientific writing (Alkaissi and McFarlane, 2023).
Several manufacturers are cautiously adopting LLMs, while seeking solutions to mitigate its associated risks. For example, (Mercedes-Benz, 2023) is utilizing AI with ChatGPT integrated through Azure OpenAI Service to enhance error identification and analysis, quality management, and process optimization in intelligent vehicle production. This AI-driven approach simplifies complex evaluations and presentations of production-related data and quality management for quality engineers through dialogue-based queries.
Also, in manufacturing, researchers have explored LLM-powered tools in knowledge-intensive scenarios. (Xia et al., 2023) demonstrated how using in-context learning and injecting task-specific knowledge into an LLM can streamline intelligent planning and control of production processes. (Kernan Freire et al., 2023) built a proof of concept for bridging knowledge gaps among workers by utilizing domain-specific texts, user requests, and query results from a knowledge graph and investigated the opportunities, risks, and user acceptance in two factories. (Wang et al., 2023) conducted a systematic test of ChatGPT’s responses to 100 questions sourced from course materials and industrial documents. They used zero-shot method and examined the responses’ correctness, relevance, clarity, and comparability. Their results suggested areas for improvement, including low scores when responding to critical analysis questions, occasional non-factual or out-of-manufacturing scope responses, and dependency on the query quality. Although (Wang et al., 2023) provides a comprehensive review of ChatGPT’s abilities to answer questions related to manufacturing, it did not include the injection of task-specific knowledge into the prompts.
To improve the performance of an LLM for domain-specific tasks, relevant context information can be automatically injected along with a question prompt. This technique, known as Retrieval Augmented Generation (RAG), involves searching a corpus for information relevant to the user’s query and inserting it into a query template before sending it to the LLM (Lewis et al., 2020). Using RAG also enables further transparency and explainability of the LLM’s response. Namely, users can check the referenced documents to verify the LLM’s response. Factories will likely have a large corpus of knowledge available in natural language, such as standard work instructions or machine manuals. Furthermore, factory workers continually add to the pool of available knowledge through (issue) reports. Until recently, these reports were considered unusable by AI natural language processing due to quality issues such as poorly structured text, inconsistent terminology, or incompleteness (Edwards et al., 2008). However, the leap in natural language understanding that LLMs, such as ChatGPT, have brought about can overcome these issues.
1.3. Evaluating Large-Language Models
Large Language Model evaluation requires the definition of evaluation criteria, metrics, and datasets associated with the system’s main tasks. There are two types of LLM evaluations: intrinsic and extrinsic evaluation. Intrinsic evaluation focuses on the internal properties of a Language Model (Wei et al., 2023). It means the patterns and language structures learned during the pre-training phase. Extrinsic evaluation focuses on the model’s performance in downstream tasks, i.e., in the execution of specific tasks that make use of the linguistic knowledge gained upstream, like code completion (Xu et al., 2022). Despite extrinsic evaluation being computationally expensive, only conducting intrinsic evaluation is not comprehensive, as it only tests the LLMs capability for memorization (Jang et al., 2022). We will focus on extrinsic evaluation as we are primarily interested in the performance of LLM-based tools for specific real-world tasks,
Extrinsic evaluation implies assessing the systems’s performance in tasks such as question answering, translation, reading comprehension, and text classification, among others (Kwon and Mihindukulasooriya, 2022). Existing benchmarks such as LAMBADA, HellaSwag, TriviaQA, BLOOM, Galactica, ClariQ and MMLU, among others, are widely reported in the literature for comparing language models. Likewise, domain-specific Benchmarks for tasks such as medical (Singhal et al., 2022), fairness evaluation (Zhang et al., 2023), finance (Xie et al., 2023), robot policies (Liang et al., 2022), and 3D printing code generation (Badini et al., 2023) can also be found.
Experts also evaluate Large-Language Models’ performance in specific downstream tasks. For example, (Lahat et al., 2023) evaluated the potential of ChatGPT for research priority identification. An expert panel was formed to review and rate the research questions generated by the system. Relevance, clarity, specificity, and originality were evaluated through a Likert scale. The interclass correlation coefficient (ICC) was also implemented to assess the inter-rater reliability.
Numerous benchmarks exist; as outlined above, they range from specific downstream tasks to general language tasks. However, to our knowledge, LLM models have not been benchmarked for answering questions in the manufacturing domain based on information retrieved from provided context material, a technique known as Retrieval Augmented Generation (Lewis et al., 2020). Material such as machine documentation, standard work instructions, or issue reports will contain domain jargon and technical information that LLMs may struggle to process. Furthermore, the text in an issue report may pose additional challenges due to poor grammar or abbreviations. Therefore, as part of this work, we benchmarked several LLM models on their ability to answer questions based on factory manuals and issue reports. Furthermore, we conducted preliminary evaluations with factory managers to assess the potential benefits, risks and challenges. The following research questions informed our study:
RQ1
What are the perceived benefits, challenges and risks of using Retrieval Augmented Generation for information retrieval and knowledge sharing for factory operators?
RQ2
How do Large Language Models (e.g., GPT-4, GPT-3.5, Llama2) compare in performance when employing Retrieval Augmented Generation over factory documentation to answer queries from operators? We consider performance as the factuality, completeness, hallucinations, and conciseness of the generated response.
2. System
We built a fully functional system to assess the potential of using LLMs for information retrieval and knowledge sharing for factory operators. Benefiting from LLMs’ in-context learning capabilities, we use this to supply an LLM with information in the form of factory manuals, and issue reports relevant to the user’s question, a technique known as Retrieval Augmented Generation (RAG), see Figure 1. As pointed out by (Wei et al., 2022), training LLMs using a prompt packed with query-related information can yield substantial performance enhancement. Users can ask questions in the chat box by typing or using voice input. The response is displayed at the top of the page, and the document chunks used for the answer can be checked at the bottom (see Figure 2).
2.0.1. Tool Dependencies
The tool was constructed utilizing two innovative technologies - Gradio and LlamaIndex. Gradio, a tool developed by Abid et al. (2019), serves as the backbone for both our front and back ends. Primarily used to simplify the development and distribution of machine learning applications, Gradio allows the quick creation of intuitive, user-friendly web interfaces for machine learning models.
Additionally, we use LlamaIndex, created by Liu (2022), for retrieving the context material in response to the user queries and handling the queries to the LLM. LlamaIndex, initially known as GPT Index, is a cutting-edge data framework designed for the efficient handling and accessibility of private or domain-specific data in LLMs applications.
2.0.2. Knowledge Base Construction
Our experiment incorporates two distinct types of domain-specific data: factory manuals and shared knowledge from factory workers. Factory manuals outline information on machine operation, safety protocols, quality assurance, and more. These resources, provided by factory management teams, initialize the knowledge base for each specific factory. The materials come in various formats, including PDF, Word, and CSV files. To ensure easy access and indexing, we leverage the capabilities of LlamaIndex.
In addition to the manuals, we integrate valuable practical knowledge from factory workers. This information is gathered from the production line, utilizing the five-why process, an iterative technique cited by (Serrat, 2017). The five-why technique probes into cause-and-effect relationships underlying specific problems by repeatedly asking ”Why?” until the root cause is revealed, typically by the fifth query. This process enables us to gather real-world issues encountered on production lines, which may not be covered in the factory manuals.
Since these documents can be extensive, integrating them directly into the LLM conversation could be challenging. To overcome this, we further segment the materials into manageable chunks, each comprising approximately 400 tokens. This method effectively incorporates the materials into the LLM prompt without compromising the conversation flow. Following the segmentation, each document chunk is processed through LlamaIndex using the OpenAI Embedding API222https://api.openai.com/v1/embeddings. Utilizing the ‘text-embedding-ada-002’ model, LlamaIndex transforms each chunk into a corresponding embedding vector. These resulting vectors are then securely stored, ready for future retrieval and use.
2.0.3. Query Construction
To retrieve the document data relevant to specific user queries, we employ the same embedding model, “text-embedding-ada-002”, to generate vector representations of these queries. By leveraging the similarity calculation algorithm provided by LlamaIndex, we can identify and retrieve the top-K most similar segmented document snippets related to the user query. This allows us to construct pertinent LLM queries. Once the snippets are retrieved, they are synthesized into the following query template based on the templates used by LlamaIndex333https://docs.llamaindex.ai:
You are an assistant that assists detergent production line operators with decision support and advice based on a knowledge base of standard operating procedures, single point lessons (SPL), etc. We have provided context information below from relevant documents and reports.
[Retrieved Document Snippets]
Given this information, please answer the following question: [Query]
If the provided context does not include relevant information to answer the question, please do not respond.
However, considering our data originates from two distinct sources – factory manuals and shared tactical knowledge – we have decided to segregate these into two separate LLM queries. This approach is designed to prevent potential user confusion from combining data from both sources into a single query.
3. Model Benchmarking
In our benchmarking experiment, we evaluated various commercial and open-source LLMs, including OpenAI’s ChatGPT (GPT-3.5 and GPT-4 from July 20th 2023), Guanaco 65B and 35B variants (Dettmers et al., 2023) based on Meta’s Llama (Large Language Model Meta AI) (Touvron et al., 2023a), Llama2 (Touvron et al., 2023b) with 70 billion parameters, and one of its derivatives, StableBeluga2 (Mahan et al., 2023).
GPT-3, or Generative Pre-trained Transformer 3, is an autoregressive language model developed by OpenAI. This third-generation model utilizes deep learning to generate human-like text from a given prompt, capable of producing various types of data such as word lists or code lines (Floridi and Chiriatti, 2020). OpenAI recently employed a Reinforcement Learning from Human Feedback (RLHF) strategy to enhance GPT-3’s reliability, usefulness, and alignment (OpenAI, 2023). This technique relies on human labellers providing desired model behavior examples and ranking responses from the GPT-3 model to specific prompts (OpenAI, 2023). Using this feedback, OpenAI fine-tunes the GPT-3 model, creating ChatGPT, a leading language model, introduced as a prototype on November 30, 2022 (OpenAI, 2023).
On the other hand, LLaMA, launched by Meta AI in February 2023, is a contrasting open-source language model (Touvron et al., 2023a). Meta AI trained a spectrum of models with parameters ranging from 7 billion to 65 billion. Like GPT-3, LLaMA also benefited from RLHF to enhance its capabilities. The LLaMA 7B model further paved the way for Alpaca’s development, a training regimen released by Taori et al. (2023). Alpaca employs a unique ”Self-Instruct” instruction tuning method, allowing it to achieve skills similar to OpenAI’s GPT-3 series (specifically the text-davinci-003 model) at a reduced cost. More recently, (Dettmers et al., 2023) released QLoRA, an efficient finetuning approach that allows 65B parameter models to be run on a single 48Gb GPU. These models, named Guanaco, are available in the following sizes: 7B, 13B, 33B, and 65B. According to the Vicuna benchmark, the 65B and 33B parameter models reached 99.3% and 97.3% ChatGPT performance, respectively. Nevertheless, like all LLaMA derivatives, they can only be used for research, not commercial applications. However, LLaMA 2 (Touvron et al., 2023b), which was released on July 18th, 2023, is available for both commercial and research use444https://about.fb.com/news/2023/07/llama-2/—last accessed January 10, 2024.
We used a web UI for LLMs555https://github.com/oobabooga/text-generation-webui/tree/main—last accessed January 10, 2024 to load and test the Guanaco model and StableBeluga2 models. The models were loaded on a pair of Nvidia A6000s with NVlink and total Video Random Access Memory (VRAM) capacity of 96 GB. The 65B model was run in 8-bit mode to fit in the available VRAM. The models were used in the chat mode, with the llama-precise preset and fixed zero seed for reproducibility. Llama2 was evaluated using the demo on huggingface666https://huggingface.co/meta-llama/Llama-2-70b-chat-hf—last accessed January 10, 2024.
To rigorously assess the models, we prepared 20 questions of varying complexity based on two types of context material: formal operating material and informal issue reports. The model prompt was constructed using the above template (2.0.3). Ultimately, the difficulty of a question is a combination of the question’s complexity and the clarity of the source material. Simple questions include retrieving a single piece of information clearly stated in the context material, for example, ”At what temperature is relubrication necessary for the OKS 4220 grease?”. Conversely, difficult questions require more reasoning or comprise multiple parts, for example, ”What should I do if the central turntable is overloaded?” which has a nuanced answer dependent on several factors not clearly articulated in the context material.
Every response is manually scored on factuality, completeness, and hallucinations. Factuality refers to aligning with the facts in the context material. Complete responses contain all the information relevant to the question in the context material. A hallucinated response appears grammatically and semantically coherent but is not based on the context material. The following scoring protocol is applied: one is awarded for a completely factual, complete, or hallucinated response. In contrast, a score of 0.5 is awarded for a slightly nonfactual, incomplete, or hallucinated response (e.g., the response includes four out of the five correct steps). Otherwise, a score of zero is awarded. Therefore, wrong answers are penalized heavily. If the model responds by saying it cannot answer the question and does not make any attempt to do so, it is scored zero for factuality and completeness, but no score is given for hallucination. As such, the final score for hallucination is calculated as follows:
As shown in Figure 3 and Table 1, GPT-4 outperforms other models regarding factuality, completeness, and lack of hallucinations but is closely followed by StableBeluga2 and GPT-3.5. The Guanaco models, based on Llama 1, perform significantly worse. The conciseness of the responses showed a similar pattern, except that StableBeluga2 produced the shortest answers (58 words), followed closely by GPT-4 (69 words).
| Model | Factuality | Completeness | Hallucinations | Words |
| GPT-4 | 97.5 | 95 | 0 | 69 |
| StableBeluga2 | 95 | 92.5 | 7.5 | 58 |
| GPT-3.5 | 90 | 90 | 5 | 89 |
| Llama2 | 77.5 | 82.5 | 13 | 128 |
| Guanaco 65B | 55 | 39.5 | 65 | 131 |
| Guanaco 33b | 27.5 | 27.5 | 65.6 | 190 |
4. User Study at the Factory
We conducted a user study on the abovementioned system we co-designed with factory employees. We recruited factory managers for the user study. The study comprised three tasks: to ask the system several questions as if they were operators, to fill in a ”yellow tag” (issue report) based on a recent issue and request a logical check, and finally, to upload new documents to the system. After each task, they were asked to provide feedback. Then, they were posed with several open questions about benefits, risks, and challenges. Finally, demographic information was requested, such as age, gender, and role.
An inductive thematic analysis of the answers to the open questions resulted in the following five themes:
Usability (5): The system is praised for its user-friendliness, speed in delivering responses, and logical control function. Users appreciate the voice command and information search capabilities (P4). However, the user interface could be improved by having attachments appear after the main text for more accessible consultation (P1).
Content (5): The system retrieves relevant information based on queries posed. Users appreciate access to relevant documents and seeing where the data is taken from. However, there’s a call for more detailed instructions to be added to the system. Also, it’s mentioned that the part of answers based on yellow tags sometimes refers to previous knowledge, causing confusion (P2).
Features (4): There are suggestions to include additional features, such as the ability to open IO, SPL, and yellow tags in .pdf format for consultation (P3). Furthermore, one participant appreciated the ability to change file names (P4). Lastly, another participant could not access the yellow tag insertion fields as the fields were in a collapsed state (P1).
Risks & Benefits (4): The system is recognized for its potential to modernize the factory and make specific tasks faster, for example, ”Greater speed in carrying out some small activities.” (P3) and ”Being able to easily find information, especially if they are in different documents.” (P4). However, concerns are raised about the safety risks if the system provides inadequate answers and the potential time-wasting if the system lacks information on specific problems (P2, P3).
Employee Acceptance & Training (2): Users hope for system acceptance among all employees. There’s also a suggestion about training and sensitization of the operators to the system’s advantages (P3). P2 implies that users may turn to a human expert if available instead of using the system as the ”[Information from the system] might be less quick, less effective, and less exhaustive than asking the expert staff present.”
5. Summary and Discussion
5.1. Model Benchmarking
GPT-4 performs best across all measures but is closely followed by StableBeluga2 and GPT-3.5. Compared to GPT-4, the cost per input token for GPT-3.5 is significantly lower777https://openai.com/pricing#language-models—last accessed January 10, 2024. The higher costs of GPT-4 are partially counteracted by the response length as GPT-3.5 (and Llama 2) tended to be wordier and include additional details that were not directly requested, whereas GPT-4 (and StableBeluga2) generated more concise responses.
The latest generation of open-source models based on Llama-2 demonstrates a clear jump forward relative to its predecessors based on Llama-1, which were more prone to hallucinations and exhibited poorer reasoning abilities over the context material. While open-source models like StableBeluga2 do not score as high as GPT-4, they ensure better data security, privacy, and customization. Unlike GPT-4 and GPT-3.5, they can be hosted locally and heavily customized. This can be a crucial consideration for companies with sensitive data or unique needs. It should be noted that fine-tuning GPT-3.5 is possible from August 22nd 2023, and it was announced that fine-tuning GPT-4 would be available in the fall of 2023 888https://openai.com/blog/gpt-3-5-turbo-fine-tuning-and-api-updates—last accessed January 10, 2024.
5.2. Perceptions at the Factory
Users appreciate the system’s functionality and see it as a tool for modernizing factory operations and speeding up operations. They are keen on improvements to be made for better user experience and utility, especially in the areas of content, feature enhancements, and user training. However, they express concerns about potential safety risks and the efficacy of information retrieval compared to consulting expert personnel. Although the tool is seen as inferior to learning from a human expert, it can help when human experts are not present or are unfamiliar with the issue at hand.
5.3. Limitations
The participant pool for the user study comprised of factory managers and was limited in size. This potentially overlooks the varied perspectives of other stakeholders, such as factory operators. To partially mitigate this we asked the participants to pose queries as if they were operators.
We used the same prompt for all LLMs; however, it is possible that some of the LLMs would perform better with a prompt template developed explicitly for it. Furthermore, we matched the settings of the LLMs (e.g., temperature) as closely as possible across all the tested models; however, the same settings across model types were not completely equivalent, and in the case of Llama 2, we did not have access to the presets as we did not host it locally.
The study’s design did not include a comprehensive real-world evaluation involving end users in their natural working environment, as this was considered too risky for our industry partner. Such an environment might present unique challenges and considerations not addressed in this research.
Our benchmarking procedure involved only 20 questions, and a singular coder assessed the responses. This introduces the potential for bias and limits the generalizability of our findings. Furthermore, the limited number of questions may not cover the full spectrum of complexities in real-world scenarios.
These limitations suggest directions for future research, where a broader participant pool, more comprehensive prompt and model customization, and real-world evaluations could be considered. Furthermore, considering the pace at which LLMs are improving, it would be desirable to (semi)automate benchmarking.
6. Conclusion
The results demonstrated GPT-4’s superior performance over other models regarding factuality, completeness, and minimal hallucinations. Interestingly, open source models like StableBeluga2 followed close behind. The user study highlighted the system’s user-friendliness, speed, and logical functionality. However, improvements in the user interface and content specificity were suggested, along with potential new features. Benefits included modernizing factory operations and speeding up certain tasks, though concerns about safety and efficiency were raised.
Acknowledgements.
This work was supported by the European Union’s Horizon 2020 research and innovation program via the project COALA “COgnitive Assisted agile manufacturing for a LAbor force supported by trustworthy Artificial Intelligence” (Grant agreement 957296).References
- (1)
- Abid et al. (2019) Abubakar Abid, Ali Abdalla, Ali Abid, Dawood Khan, Abdulrahman Alfozan, and James Zou. 2019. Gradio: Hassle-free sharing and testing of ML models in the wild. https://doi.org/10.48550/arXiv.1906.02569
- Alkaissi and McFarlane (2023) Hussam Alkaissi and Samy I. McFarlane. 2023. Artificial Hallucinations in ChatGPT: Implications in Scientific Writing. Cureus 15, 2 (2023), e35179. https://doi.org/10.7759/cureus.35179
- Alves et al. (2023) Joel Alves, Tânia M. Lima, and Pedro D. Gaspar. 2023. Is Industry 5.0 a Human-Centred Approach? A Systematic Review. Processes 11, 1 (2023). https://doi.org/10.3390/pr11010193
- Badini et al. (2023) Silvia Badini, Stefano Regondi, Emanuele Frontoni, and Raffaele Pugliese. 2023. Assessing the capabilities of ChatGPT to improve additive manufacturing troubleshooting. Advanced Industrial and Engineering Polymer Research (3 2023). https://doi.org/10.1016/J.AIEPR.2023.03.003
- Bang et al. (2023) Yejin Bang, Samuel Cahyawijaya, Nayeon Lee, Wenliang Dai, Dan Su, Bryan Wilie, Holy Lovenia, Ziwei Ji, Tiezheng Yu, Willy Chung, Quyet V. Do, Yan Xu, and Pascale Fung. 2023. A Multitask, Multilingual, Multimodal Evaluation of ChatGPT on Reasoning, Hallucination, and Interactivity. arXiv:2302.04023 [cs.CL]
- Brown et al. (2020) Tom Brown, Benjamin Mann, Nick Ryder, Melanie Subbiah, Jared D Kaplan, Prafulla Dhariwal, Arvind Neelakantan, Pranav Shyam, Girish Sastry, Amanda Askell, Sandhini Agarwal, Ariel Herbert-Voss, Gretchen Krueger, Tom Henighan, Rewon Child, Aditya Ramesh, Daniel Ziegler, Jeffrey Wu, Clemens Winter, Chris Hesse, Mark Chen, Eric Sigler, Mateusz Litwin, Scott Gray, Benjamin Chess, Jack Clark, Christopher Berner, Sam McCandlish, Alec Radford, Ilya Sutskever, and Dario Amodei. 2020. Language Models are Few-Shot Learners. In Advances in Neural Information Processing Systems, H. Larochelle, M. Ranzato, R. Hadsell, M.F. Balcan, and H. Lin (Eds.), Vol. 33. Curran Associates, Inc., 1877–1901. https://proceedings.neurips.cc/paper_files/paper/2020/file/1457c0d6bfcb4967418bfb8ac142f64a-Paper.pdf
- Brückner et al. (2023) Anja Brückner, Philipp Hein, Franziska Hein-Pensel, Jasmin Mayan, and Mandy Wölke. 2023. Human-Centered HCI Practices Leading the Path to Industry 5.0: A Systematic Literature Review. In HCI International 2023 Posters, Constantine Stephanidis, Margherita Antona, Stavroula Ntoa, and Gavriel Salvendy (Eds.). Springer Nature Switzerland, Cham, 3–15.
- Dettmers et al. (2023) Tim Dettmers, Artidoro Pagnoni, Ari Holtzman, and Luke Zettlemoyer. 2023. QLoRA: Efficient Finetuning of Quantized LLMs. arXiv:2305.14314 [cs.LG]
- Edwards et al. (2008) Brett Edwards, Michael Zatorsky, and Richi Nayak. 2008. Clustering and classification of maintenance logs using text data mining. Volume 87-Data Mining and Analytics 2008 (2008), 193–199.
- Floridi and Chiriatti (2020) Luciano Floridi and Massimo Chiriatti. 2020. GPT-3: Its nature, scope, limits, and consequences. Minds and Machines 30, 4 (2020), 681–694.
- Gao et al. (2021) Tianyu Gao, Adam Fisch, and Danqi Chen. 2021. Making Pre-trained Language Models Better Few-shot Learners. In Proceedings of the 59th Annual Meeting of the Association for Computational Linguistics and the 11th International Joint Conference on Natural Language Processing (Volume 1: Long Papers). Association for Computational Linguistics, Online, 3816–3830. https://doi.org/10.18653/v1/2021.acl-long.295
- Jang et al. (2022) Joel Jang, Seonghyeon Ye, Changho Lee, Sohee Yang, Joongbo Shin, Janghoon Han, Gyeonghun Kim, and Minjoon Seo. 2022. TemporalWiki: A Lifelong Benchmark for Training and Evaluating Ever-Evolving Language Models. (2022). https://commons.wikimedia.org/
- Jawahar et al. (2019) Ganesh Jawahar, Benoît Sagot, and Djamé Seddah. 2019. What Does BERT Learn about the Structure of Language?. In Proceedings of the 57th Annual Meeting of the Association for Computational Linguistics. Association for Computational Linguistics, Florence, Italy, 3651–3657. https://doi.org/10.18653/v1/P19-1356
- Jeblick et al. (2022) Katharina Jeblick, Balthasar Schachtner, Jakob Dexl, Andreas Mittermeier, Anna Theresa Stüber, Johanna Topalis, Tobias Weber, Philipp Wesp, Bastian Sabel, Jens Ricke, and Michael Ingrisch. 2022. ChatGPT Makes Medicine Easy to Swallow: An Exploratory Case Study on Simplified Radiology Reports. arXiv:2212.14882 [cs.CL]
- Kernan Freire et al. (2023) Samuel Kernan Freire, Mina Foosherian, Chaofan Wang, and Evangelos Niforatos. 2023. Harnessing Large Language Models for Cognitive Assistants in Factories. In Proceedings of the 5th International Conference on Conversational User Interfaces (Eindhoven, Netherlands) (CUI ’23). Association for Computing Machinery, New York, NY, USA, Article 44, 6 pages. https://doi.org/10.1145/3571884.3604313
- Kwon and Mihindukulasooriya (2022) Bum Chul Kwon and Nandana Mihindukulasooriya. 2022. An Empirical Study on Pseudo-log-likelihood Bias Measures for Masked Language Models Using Paraphrased Sentences. TrustNLP 2022 - 2nd Workshop on Trustworthy Natural Language Processing, Proceedings of the Workshop (2022), 74–79. https://doi.org/10.18653/V1/2022.TRUSTNLP-1.7
- Lahat et al. (2023) Adi Lahat, Eyal Shachar, Benjamin Avidan, Zina Shatz, Benjamin S. Glicksberg, and Eyal Klang. 2023. Evaluating the use of large language model in identifying top research questions in gastroenterology. Scientific Reports 2023 13:1 13 (3 2023), 1–6. Issue 1. https://doi.org/10.1038/s41598-023-31412-2
- Lewis et al. (2020) Patrick Lewis, Ethan Perez, Aleksandra Piktus, Fabio Petroni, Vladimir Karpukhin, Naman Goyal, Heinrich Küttler, Mike Lewis, Wen-tau Yih, Tim Rocktäschel, Sebastian Riedel, and Douwe Kiela. 2020. Retrieval-Augmented Generation for Knowledge-Intensive NLP Tasks. In Proceedings of the 34th International Conference on Neural Information Processing Systems (Vancouver, BC, Canada) (NIPS’20). Curran Associates Inc., Red Hook, NY, USA, Article 793, 16 pages.
- Liang et al. (2022) Jacky Liang, Wenlong Huang, Fei Xia, Peng Xu, Karol Hausman, Brian Ichter, Pete Florence, and Andy Zeng. 2022. Code as Policies: Language Model Programs for Embodied Control. (9 2022). https://arxiv.org/abs/2209.07753v4
- Liu (2022) Jerry Liu. 2022. LlamaIndex. https://doi.org/10.5281/zenodo.1234
- Maddikunta et al. (2022) Praveen Kumar Reddy Maddikunta, Quoc-Viet Pham, Prabadevi B, N Deepa, Kapal Dev, Thippa Reddy Gadekallu, Rukhsana Ruby, and Madhusanka Liyanage. 2022. Industry 5.0: A survey on enabling technologies and potential applications. Journal of Industrial Information Integration 26 (2022), 100257. https://doi.org/10.1016/j.jii.2021.100257
- Mahan et al. (2023) Dakota Mahan, Ryan Carlow, Louis Castricato, Nathan Cooper, and Christian Laforte. 2023. Stable Beluga models. [https://huggingface.co/stabilityai/StableBeluga2](https://huggingface.co/stabilityai/StableBeluga2)
- Mercedes-Benz (2023) Mercedes-Benz. 2023. Benz tests chatgpt in intelligent vehicle production.: Mercedes-Benz Group. https://group.mercedes-benz.com/innovation/digitalisation/industry-4-0/chatgpt-in-vehicle-production.html
- Nov et al. (2023) Oded Nov, Nina Singh, and Devin Mann. 2023. Putting ChatGPT’s Medical Advice to the (Turing) Test. arXiv:2301.10035 [cs.HC]
- OpenAI (2023) OpenAI. 2023. ChatGPT: Optimizing Language Models for Dialogue. https://openai.com/blog/chatgpt/.
- Semnani et al. (2023) Sina J. Semnani, Violet Z. Yao, Heidi C. Zhang, and Monica S. Lam. 2023. WikiChat: A Few-Shot LLM-Based Chatbot Grounded with Wikipedia. arXiv:2305.14292 [cs.CL]
- Serrat (2017) Olivier Serrat. 2017. The five whys technique. Knowledge solutions: Tools, methods, and approaches to drive organizational performance (2017), 307–310.
- Singhal et al. (2022) Karan Singhal, Shekoofeh Azizi, Tao Tu, S. Sara Mahdavi, Jason Wei, Hyung Won Chung, Nathan Scales, Ajay Tanwani, Heather Cole-Lewis, Stephen Pfohl, Perry Payne, Martin Seneviratne, Paul Gamble, Chris Kelly, Nathaneal Scharli, Aakanksha Chowdhery, Philip Mansfield, Blaise Aguera y Arcas, Dale Webster, Greg S. Corrado, Yossi Matias, Katherine Chou, Juraj Gottweis, Nenad Tomasev, Yun Liu, Alvin Rajkomar, Joelle Barral, Christopher Semturs, Alan Karthikesalingam, and Vivek Natarajan. 2022. Large Language Models Encode Clinical Knowledge. arXiv:2212.13138 [cs.CL]
- Tang et al. (2023) Ruixiang Tang, Xiaotian Han, Xiaoqian Jiang, and Xia Hu. 2023. Does Synthetic Data Generation of LLMs Help Clinical Text Mining? arXiv:2303.04360 [cs.CL]
- Taori et al. (2023) Rohan Taori, Ishaan Gulrajani, Tianyi Zhang, Yann Dubois, Xuechen Li, Carlos Guestrin, Percy Liang, and Tatsunori B. Hashimoto. 2023. Stanford Alpaca: An Instruction-following LLaMA model. https://github.com/tatsu-lab/stanford_alpaca.
- Touvron et al. (2023a) Hugo Touvron, Louis Martin, Kevin Stone, Peter Albert, Amjad Almahairi, Yasmine Babaei, Nikolay Bashlykov, Soumya Batra, Prajjwal Bhargava, Shruti Bhosale, Dan Bikel, Lukas Blecher, Cristian Canton Ferrer, Moya Chen, Guillem Cucurull, David Esiobu, Jude Fernandes, Jeremy Fu, Wenyin Fu, Brian Fuller, Cynthia Gao, Vedanuj Goswami, Naman Goyal, Anthony Hartshorn, Saghar Hosseini, Rui Hou, Hakan Inan, Marcin Kardas, Viktor Kerkez, Madian Khabsa, Isabel Kloumann, Artem Korenev, Punit Singh Koura, Marie-Anne Lachaux, Thibaut Lavril, Jenya Lee, Diana Liskovich, Yinghai Lu, Yuning Mao, Xavier Martinet, Todor Mihaylov, Pushkar Mishra, Igor Molybog, Yixin Nie, Andrew Poulton, Jeremy Reizenstein, Rashi Rungta, Kalyan Saladi, Alan Schelten, Ruan Silva, Eric Michael Smith, Ranjan Subramanian, Xiaoqing Ellen Tan, Binh Tang, Ross Taylor, Adina Williams, Jian Xiang Kuan, Puxin Xu, Zheng Yan, Iliyan Zarov, Yuchen Zhang, Angela Fan, Melanie Kambadur, Sharan Narang, Aurelien Rodriguez, Robert Stojnic, Sergey Edunov, and Thomas Scialom. 2023a. Llama 2: Open Foundation and Fine-Tuned Chat Models. arXiv:2307.09288 [cs.CL]
- Touvron et al. (2023b) Hugo Touvron, Louis Martin, Kevin Stone, Peter Albert, Amjad Almahairi, Yasmine Babaei, Nikolay Bashlykov, Soumya Batra, Prajjwal Bhargava, Shruti Bhosale, Dan Bikel, Lukas Blecher, Cristian Canton Ferrer, Moya Chen, Guillem Cucurull, David Esiobu, Jude Fernandes, Jeremy Fu, Wenyin Fu, Brian Fuller, Cynthia Gao, Vedanuj Goswami, Naman Goyal, Anthony Hartshorn, Saghar Hosseini, Rui Hou, Hakan Inan, Marcin Kardas, Viktor Kerkez, Madian Khabsa, Isabel Kloumann, Artem Korenev, Punit Singh Koura, Marie-Anne Lachaux, Thibaut Lavril, Jenya Lee, Diana Liskovich, Yinghai Lu, Yuning Mao, Xavier Martinet, Todor Mihaylov, Pushkar Mishra, Igor Molybog, Yixin Nie, Andrew Poulton, Jeremy Reizenstein, Rashi Rungta, Kalyan Saladi, Alan Schelten, Ruan Silva, Eric Michael Smith, Ranjan Subramanian, Xiaoqing Ellen Tan, Binh Tang, Ross Taylor, Adina Williams, Jian Xiang Kuan, Puxin Xu, Zheng Yan, Iliyan Zarov, Yuchen Zhang, Angela Fan, Melanie Kambadur, Sharan Narang, Aurelien Rodriguez, Robert Stojnic, Sergey Edunov, and Thomas Scialom. 2023b. Llama 2: Open Foundation and Fine-Tuned Chat Models. arXiv:2307.09288 [cs.CL]
- Trautmann et al. (2022) Dietrich Trautmann, Alina Petrova, and Frank Schilder. 2022. Legal Prompt Engineering for Multilingual Legal Judgement Prediction. arXiv:2212.02199 [cs.CL]
- Wang et al. (2023) Xingzhi Wang, Nabil Anwer, Yun Dai, and Ang Liu. 2023. ChatGPT for design, manufacturing, and education. Procedia CIRP 119 (2023), 7–14. https://doi.org/10.1016/j.procir.2023.04.001 The 33rd CIRP Design Conference.
- Wei et al. (2023) Chengwei Wei, Yun-Cheng Wang, Bin Wang, and C. C. Jay Kuo. 2023. An Overview on Language Models: Recent Developments and Outlook. (3 2023). https://arxiv.org/abs/2303.05759v1
- Wei et al. (2022) Jason Wei, Yi Tay, Rishi Bommasani, Colin Raffel, Barret Zoph, Sebastian Borgeaud, Dani Yogatama, Maarten Bosma, Denny Zhou, Donald Metzler, Ed H. Chi, Tatsunori Hashimoto, Oriol Vinyals, Percy Liang, Jeff Dean, and William Fedus. 2022. Emergent Abilities of Large Language Models. arXiv:2206.07682 [cs.CL]
- Wellsandt et al. (2021) Stefan Wellsandt, Karl Hribernik, and Klaus-Dieter Thoben. 2021. Anatomy of a Digital Assistant. In Advances in Production Management Systems. Artificial Intelligence for Sustainable and Resilient Production Systems, Alexandre Dolgui, Alain Bernard, David Lemoine, Gregor von Cieminski, and David Romero (Eds.). Springer International Publishing, Cham, 321–330.
- Xia et al. (2023) Yuchen Xia, Manthan Shenoy, Nasser Jazdi, and Michael Weyrich. 2023. Towards autonomous system: flexible modular production system enhanced with large language model agents. arXiv:2304.14721 [cs.RO]
- Xie et al. (2023) Qianqian Xie, Weiguang Han, Xiao Zhang, Yanzhao Lai, Min Peng, Alejandro Lopez-Lira, and Jimin Huang. 2023. PIXIU: A Large Language Model, Instruction Data and Evaluation Benchmark for Finance. (6 2023). https://arxiv.org/abs/2306.05443v1
- Xu et al. (2022) Frank F Xu, Uri Alon, Graham Neubig, Vincent Josua Hellendoorn, and Vincent Josua Hel. 2022. A systematic evaluation of large language models of code. (6 2022), 1–10. https://doi.org/10.1145/3520312.3534862
- Xu et al. (2021) Xun Xu, Yuqian Lu, Birgit Vogel-Heuser, and Lihui Wang. 2021. Industry 4.0 and Industry 5.0—Inception, conception and perception. Journal of Manufacturing Systems 61 (2021), 530–535. https://doi.org/10.1016/j.jmsy.2021.10.006
- Zhang et al. (2023) Jianzhang Zhang, Yiyang Chen, Nan Niu, and Chuang Liu. 2023. A Preliminary Evaluation of ChatGPT in Requirements Information Retrieval. (4 2023). https://arxiv.org/abs/2304.12562v1
- Zhao et al. (2023) Wayne Xin Zhao, Kun Zhou, Junyi Li, Tianyi Tang, Xiaolei Wang, Yupeng Hou, Yingqian Min, Beichen Zhang, Junjie Zhang, Zican Dong, Yifan Du, Chen Yang, Yushuo Chen, Zhipeng Chen, Jinhao Jiang, Ruiyang Ren, Yifan Li, Xinyu Tang, Zikang Liu, Peiyu Liu, Jian-Yun Nie, and Ji-Rong Wen. 2023. A Survey of Large Language Models. arXiv:2303.18223 [cs.CL]

### 26 Feb 2024 v2

@@
 Abstract.
-Managing knowledge efficiently is crucial for organizational success.
-In manufacturing, operating factories has become increasing knowledge-intensive putting strain on the factory’s capacity to train and support new operators.
-In this paper, we introduce a Large Language Model (LLM)-based system designed to use the extensive knowledge contained in factory documentation.
+Recent advances in natural language processing enable more intelligent ways to support knowledge sharing in factories.
+In manufacturing, operating production lines has become increasingly knowledge-intensive, putting strain on a factory’s capacity to train and support new operators.
+This paper introduces a Large Language Model (LLM)-based system designed to retrieve information from the extensive knowledge contained in factory documentation and knowledge shared by expert operators.
 The system aims to efficiently answer queries from operators and facilitate the sharing of new knowledge.
-To assess its effectiveness, we conducted an evaluation in a factory setting.
-The results of this evaluation demonstrated the system’s benefits; namely, in enabling quicker information retrieval and more efficient resolution of issues.
+We conducted a user study at a factory to assess its potential impact and adoption, eliciting several perceived benefits, namely, enabling quicker information retrieval and more efficient resolution of issues.
 However, the study also highlighted a preference for learning from a human expert when such an option is available.
-Furthermore, we benchmarked several closed and open-sourced LLMs for this system.
-GPT-4 consistently outperformed its counterparts, with open-source models like StableBeluga2 trailing closely, presenting an attractive option given its data privacy and customization benefits.
-Overall, this work offers preliminary insights for factories considering using LLM-tools for knowledge management.
+Furthermore, we benchmarked several commercial and open-sourced LLMs for this system.
+The current state-of-the-art model, GPT-4, consistently outperformed its counterparts, with open-source models trailing closely, presenting an attractive option given their data privacy and customization benefits.
+In summary, this work offers preliminary insights and a system design for factories considering using LLM tools for knowledge management.
 1. Introduction
-Managing knowledge has become invaluable in today’s manufacturing industry, and technology plays an ever-increasing role.
-Factories already generated a wealth of knowledge in the form of issue reports and machine documentation; however, up till now, these sources have been challenging to harness using AI.
-With the advent of Large Language Models, such as GPT-4, the potential to use technology to support operators in knowledge-intensive tasks has leaped forward.
-However, using LLMs comes with many (unknown) practical and socio-technical risks and challenges.
-Furthermore, a generic LLM has limited use in a factory, where specific knowledge is needed.
-As such, an LLM-powered tool for manufacturing would require fine-tuning the model or including context information in its prompt (e.g., sections from a machine manual).
-We built a fully functional LLM-powered tool for this work that finds relevant sections from uploaded documents and issue reports to answer operator queries.
-We conducted a user study at a factory to evaluate the usability and potential of the system.
-Furthermore, we benchmarked several LLMs for the system, including closed and open source options.
-LLMs are typically benchmarked by the factuality of their response to general knowledge questions, resulting in LLM leaderboards111https://huggingface.co/spaces/HuggingFaceH4/open_llm_leaderboard.
-However, the ability to answer general knowledge questions is less relevant to a manufacturing environment where very specific, dynamic knowledge is required.
-Our findings could help decide how factories use this technology in the future.
-1.1. Industry 5.0
-Industry 5.0, the latest phase of industrial development, places human beings at the forefront of manufacturing processes, emphasizing their skills, creativity, and problem-solving abilities (Alves et al., 2023).
-This human-centered approach addresses the dehumanization of work, which has been a challenge in previous industrial revolutions.
-Industry 5.0 seeks to create a more fulfilling and productive work environment by empowering the workforce.
-One of the fundamental problems that Industry 5.0 addresses is the alienation and dissatisfaction experienced by workers as a result of automation and machine dominance (Alves et al., 2023).
-The rapid advancement of technology has often led to a perception that humans are mere operators, overshadowed by the efficiency and precision of machines.
-However, in Industry 5.0, humans regain their significance in manufacturing.
-By integrating advanced technologies such as robotics and artificial intelligence, Industry 5.0 enables humans to collaborate with machines, leveraging their complementary strengths.
-This collaboration enhances productivity and efficiency and acknowledges the value of human contributions.
-Human-centered manufacturing in Industry 5.0 focuses on providing a work environment that nurtures individuals’ creativity and problem-solving capabilities (Maddikunta et al., 2022).
+Human-centric manufacturing seeks to support human operators with technology, aiming to enhance creativity, human well-being, problem-solving abilities, and overall productivity within factories (Alves et al., 2023; May et al., 2015; Fantini et al., 2020).
+However, a significant challenge persists in effectively managing and utilizing the vast knowledge generated within these manufacturing environments, such as issue reports and machine documentation (Man, 2014).
+This knowledge is crucial for optimizing operations, yet it remains largely untapped due to the difficulties in processing and interpreting the disconnected, sometimes unstructured, technical information it contains (Leoni et al., 2022; Edwards et al., 2008).
+Traditionally, leveraging this knowledge has been cumbersome, with operators choosing to use personal smartphones over official procedures (Richter et al., 2019) and AI unable to handle the complexity of the data (Edwards et al., 2008).
+However, recent Large Language Models (LLMs) like GPT-4 show promise in addressing these challenges.
+LLMs can effectively interpret, summarize, and retrieve information from vast text-based datasets (Lewis et al., 2020) while concurrently aiding the capture of new knowledge (Kernan Freire et al., 2023b).
+These capabilities could significantly support operators in knowledge-intensive tasks, making it easier to access relevant information, share new knowledge, and make informed decisions rapidly.
+While LLMs offer promising capabilities, their application in manufacturing is not straightforward.
+The specific, dynamic knowledge required in this domain poses unique challenges (Feng et al., 2017).
+For instance, a foundational LLM may have limited utility in a factory setting without significant customization, such as fine-tuning or incorporating specific context information into its prompts (Wang et al., 2023b).
+Additionally, the practical and socio-technical risks and challenges of deploying LLMs in such environments remain largely unexplored — factors key to human-centered AI (Shneiderman, 2022).
+Concerns include the accuracy of the information provided, the potential for ”hallucinated” answers (Zuccon et al., 2023), and the need for systems that can adapt to the highly specialized and evolving knowledge base of a specific manufacturing setting (Feng et al., 2017).
+In response to these challenges, we developed an LLM-powered tool to leverage factory documents and issue analysis reports to answer operators’ queries.
+Furthermore, the tool facilitates the analysis and reporting of new issues.
+This tool demonstrates the feasibility of using LLMs to enhance knowledge management in manufacturing settings.
+To understand its effectiveness and potential, we conducted a user study in a factory environment, evaluating the system’s usability, user perceptions, adoption, and impact on factory operations.
+Our approach also addresses the lack of specific benchmarks for evaluating LLMs in manufacturing.
+We benchmarked several LLMs, including both closed and open-source models, recognizing that the standard benchmarks111https://huggingface.co/spaces/HuggingFaceH4/open_llm_leaderboard—last accessed February 26, 2024 primarily focus on general knowledge and reasoning.
+As such, they may not adequately reflect the challenges of understanding manufacturing-specific terminology and concepts.
+This benchmarking focused on their ability to utilize factory-specific documents and unstructured issue reports to provide factual and complete answers to operators’ queries.
+2. Background
+In this section, we address the topic of industry 5.0, LLM-powered tools for knowledge management, benchmarking LLMs, and the research questions informing this work.
+2.1. Human-centered Manufacturing
+Industry 5.0, the latest phase of industrial development, places human beings at the forefront of manufacturing processes, emphasizing their skills, creativity, and problem-solving abilities (Xu et al., 2021a; Maddikunta et al., 2022b; Alves et al., 2023).
+Human-centered manufacturing in Industry 5.0 focuses on providing a work environment that nurtures individuals’ creativity and problem-solving capabilities (Maddikunta et al., 2022a).
 It encourages workers to think critically, innovate, and continuously learn.
 With machines handling repetitive and mundane tasks, human workers can dedicate their time and energy to more complex and intellectually stimulating activities.
-This shift could enhance job satisfaction and promote personal and professional growth, as workers could acquire new skills and engage in higher-level decision-making (Alves et al., 2023), (Xu et al., 2021).
-Emphasis on human-machine collaboration and the continuous emergence and refinement of technology increase the need for adequate human-computer interaction (Brückner et al., 2023).
+This shift could enhance job satisfaction and promote personal and professional growth, as workers could acquire new skills and engage in higher-level decision-making (Alves et al., 2023; Xu et al., 2021b).
+Emphasis on human-machine collaboration and the continuous emergence and refinement of technology increases the need for adequate human-computer interaction (Brückner et al., 2023).
 One of the approaches to address this topic is using conversational AI to assist humans in manufacturing (Wellsandt et al., 2021).
-1.2. Large-Language Model-Powered Tools for Knowledge-Intensive Scenarios
-The pre-training of LLMs on large-scale, mixed-source corpora allows them to capture extensive knowledge (Zhao et al., 2023), and demonstrate proficiency in interpreting complex information (Jawahar et al., 2019), generating general reasoning and insights (Wei et al., 2022), and aiding knowledge-intensive decision-making in various contexts.
-Consequently, researchers have been exploring LLMs adoption in both general and domain-specific tasks.
-However, despite their advantages, the responses generated by LLMs in the mentioned scenarios may encounter two potential issues: (1) outdated information originating from the model’s pre-training date, and (2) inaccuracies in factual representation, known as “hallucination”(Bang et al., 2023; Zhao et al., 2023).
-To address these challenges and leverage the capabilities of LLMs in domain-specific knowledge-intensive tasks, chain-of-thought and few-shot prompting (Brown et al., 2020; Gao et al., 2021) techniques are utilized.
-This section explores LLM-based tools designed to acquire information and support decision making in both non-manufacturing and manufacturing settings.
-Using few-shot prompting to retrieve information across diverse topics, (Semnani et al., 2023) introduced an open-domain LLM-powered chatbot called WikiChat.
+2.2. LLM-powered Knowledge Management Tools
+Training Large Language Models (LLMs) on numerous, diverse texts results in the embedding of extensive knowledge (Zhao et al., 2023).
+LLMs can also adeptly interpret complex information (Jawahar et al., 2019), general reasoning (Wei et al., 2022a), and aiding knowledge-intensive decision-making.
+Consequently, researchers have been exploring applying LLM-powered tools in domain-specific tasks (Xie et al., 2023b; Wen et al., 2023; Zhang et al., 2023b).
+Despite their potential benefits, the responses generated by LLMs may have two potential issues: (1) outdated information originating from the model’s training date, and (2) inaccuracies in factual representation, known as “hallucinations” (Bang et al., 2023; Zhao et al., 2023).
+To address these challenges and leverage the capabilities of LLMs in domain-specific knowledge-intensive tasks, several techniques can be used, such as chain-of-thought (Wei et al., 2022b), few-shot prompting (Brown et al., 2020; Gao et al., 2021), and retrieval augmented generation (Lewis et al., 2020).
+Using few-shot prompting to retrieve information across diverse topics, Semnani et al.
+(2023) introduced an open-domain LLM-powered chatbot called WikiChat.
 WikiChat utilizes a 7-stage pipeline of few-shot prompted LLM that suggests facts verified against Wikipedia, retrieves additional up-to-date information, and generates coherent responses.
-They used a hybrid human-and-LLM method to evaluate the chatbot on different topics for factuality, alignment with real-worth truths and verifiable facts, and conversationality, a compound metric that scores how informational, natural, non-repetitive, and temporally correct the response is.
+They used a hybrid human-and-LLM method to evaluate the chatbot on different topics for factuality, alignment with real-worth truths and verifiable facts, and conversationality.
+This compound metric scores how informational, natural, non-repetitive, and temporally correct the response is.
 Their solution significantly outperforms GPT-3.5 in factuality, with an average improvement of 24.4% while staying on par in conversationality.
 Others have explored the capabilities of LLMs in domain-specific tasks such as extracting structured data from unstructured healthcare texts (Tang et al., 2023), providing medical advice (Nov et al., 2023), simplifying radiology reports (Jeblick et al., 2022), Legal Judgement Prediction from multilingual legal documents (Trautmann et al., 2022), and scientific writing (Alkaissi and McFarlane, 2023).
-Several manufacturers are cautiously adopting LLMs, while seeking solutions to mitigate its associated risks.
-For example, (Mercedes-Benz, 2023) is utilizing AI with ChatGPT integrated through Azure OpenAI Service to enhance error identification and analysis, quality management, and process optimization in intelligent vehicle production.
-This AI-driven approach simplifies complex evaluations and presentations of production-related data and quality management for quality engineers through dialogue-based queries.
-Also, in manufacturing, researchers have explored LLM-powered tools in knowledge-intensive scenarios.
-(Xia et al., 2023) demonstrated how using in-context learning and injecting task-specific knowledge into an LLM can streamline intelligent planning and control of production processes.
-(Kernan Freire et al., 2023) built a proof of concept for bridging knowledge gaps among workers by utilizing domain-specific texts, user requests, and query results from a knowledge graph and investigated the opportunities, risks, and user acceptance in two factories.
-(Wang et al., 2023) conducted a systematic test of ChatGPT’s responses to 100 questions sourced from course materials and industrial documents.
-They used zero-shot method and examined the responses’ correctness, relevance, clarity, and comparability.
-Their results suggested areas for improvement, including low scores when responding to critical analysis questions, occasional non-factual or out-of-manufacturing scope responses, and dependency on the query quality.
-Although (Wang et al., 2023) provides a comprehensive review of ChatGPT’s abilities to answer questions related to manufacturing, it did not include the injection of task-specific knowledge into the prompts.
+Several manufacturers are cautiously adopting LLMs, while seeking solutions to mitigate their associated risks.
+For example, Mercedes-Benz (2023) used AI with ChatGPT integrated through Azure OpenAI Service to enhance quality management and process optimization in vehicle production.
+This AI-driven approach simplifies complex evaluations for quality engineers through dialogue-based queries.
+Xia et al.
+(2023) demonstrated how using in-context learning and injecting task-specific knowledge into an LLM can streamline intelligent planning and control of production processes.
+Kernan Freire et al.
+(2023a) built a proof of concept for bridging knowledge gaps among workers by utilizing domain-specific texts and knowledge graphs.
+Wang et al.
+(2023a) conducted a systematic test of ChatGPT’s responses to 100 questions from course materials and industrial documents.
+They used a zero-shot method and examined the responses’ correctness, relevance, clarity, and comparability.
+Their results suggested areas for improvement, including low scores when responding to critical analysis questions, occasional non-factual or out-of-manufacturing scope responses, and dependency on query quality.
+Although (Wang et al., 2023a) provides a comprehensive review of ChatGPT’s abilities to answer questions related to manufacturing; it did not include the injection of task-specific knowledge into the prompts.
 To improve the performance of an LLM for domain-specific tasks, relevant context information can be automatically injected along with a question prompt.
@@
 However, the leap in natural language understanding that LLMs, such as ChatGPT, have brought about can overcome these issues.
-1.3. Evaluating Large-Language Models
+2.3. Evaluating LLMs
 Large Language Model evaluation requires the definition of evaluation criteria, metrics, and datasets associated with the system’s main tasks.
@@
 Despite extrinsic evaluation being computationally expensive, only conducting intrinsic evaluation is not comprehensive, as it only tests the LLMs capability for memorization (Jang et al., 2022).
-We will focus on extrinsic evaluation as we are primarily interested in the performance of LLM-based tools for specific real-world tasks,
+Here, we focus on extrinsic evaluation as we are primarily interested in the performance of LLM-based tools for specific real-world tasks.
 Extrinsic evaluation implies assessing the systems’s performance in tasks such as question answering, translation, reading comprehension, and text classification, among others (Kwon and Mihindukulasooriya, 2022).
 Existing benchmarks such as LAMBADA, HellaSwag, TriviaQA, BLOOM, Galactica, ClariQ and MMLU, among others, are widely reported in the literature for comparing language models.
-Likewise, domain-specific Benchmarks for tasks such as medical (Singhal et al., 2022), fairness evaluation (Zhang et al., 2023), finance (Xie et al., 2023), robot policies (Liang et al., 2022), and 3D printing code generation (Badini et al., 2023) can also be found.
-Experts also evaluate Large-Language Models’ performance in specific downstream tasks.
-For example, (Lahat et al., 2023) evaluated the potential of ChatGPT for research priority identification.
-An expert panel was formed to review and rate the research questions generated by the system.
-Relevance, clarity, specificity, and originality were evaluated through a Likert scale.
-The interclass correlation coefficient (ICC) was also implemented to assess the inter-rater reliability.
-Numerous benchmarks exist; as outlined above, they range from specific downstream tasks to general language tasks.
-However, to our knowledge, LLM models have not been benchmarked for answering questions in the manufacturing domain based on information retrieved from provided context material, a technique known as Retrieval Augmented Generation (Lewis et al., 2020).
+Likewise, domain-specific Benchmarks for tasks such as medical (Singhal et al., 2022), fairness evaluation (Zhang et al., 2023a), finance (Xie et al., 2023a), robot policies (Liang et al., 2022), and 3D printing code generation (Badini et al., 2023) can also be found.
+Experts also evaluate the performance of large-language models (LLMs) in specific downstream tasks, such as using physicians to evaluate the output of medical specific LLMs (Singhal et al., 2022).
+LLM benchmarks range from specific downstream tasks to general language tasks.
+However, to our knowledge, LLMs have not been benchmarked for answering questions in the manufacturing domain based on context material, a technique known as Retrieval Augmented Generation (Lewis et al., 2020).
 Material such as machine documentation, standard work instructions, or issue reports will contain domain jargon and technical information that LLMs may struggle to process.
-Furthermore, the text in an issue report may pose additional challenges due to poor grammar or abbreviations.
-Therefore, as part of this work, we benchmarked several LLM models on their ability to answer questions based on factory manuals and issue reports.
-Furthermore, we conducted preliminary evaluations with factory managers to assess the potential benefits, risks and challenges.
+Furthermore, the text in an issue report may pose additional challenges due to abbreviations, poor grammar, and formatting (Edwards et al., 2008).
+Therefore, as part of this work, we benchmarked several LLMs on their ability to answer questions based on factory manuals and unstructured issue reports.
+Furthermore, we conducted a user study with factory operators and managers to assess the potential benefits, risks and challenges.
 The following research questions informed our study:
-RQ1
-What are the perceived benefits, challenges and risks of using Retrieval Augmented Generation for information retrieval and knowledge sharing for factory operators?
-RQ2
-How do Large Language Models (e.g., GPT-4, GPT-3.5, Llama2) compare in performance when employing Retrieval Augmented Generation over factory documentation to answer queries from operators?
+What are the perceived benefits, challenges, and risks of using Large Language Models for information retrieval and knowledge sharing for factory operators?
+How do Large Language Models compare in performance when answering factory operators’ queries based on factory documentation and unstructured issue reports?
 We consider performance as the factuality, completeness, hallucinations, and conciseness of the generated response.
@@
 We built a fully functional system to assess the potential of using LLMs for information retrieval and knowledge sharing for factory operators.
-Benefiting from LLMs’ in-context learning capabilities, we use this to supply an LLM with information in the form of factory manuals, and issue reports relevant to the user’s question, a technique known as Retrieval Augmented Generation (RAG), see Figure 1.
-As pointed out by (Wei et al., 2022), training LLMs using a prompt packed with query-related information can yield substantial performance enhancement.
+Benefiting from LLMs’ in-context learning capabilities, we use this to supply an LLM with information in the form of factory manuals, and issue reports relevant to the user’s question, a technique known as Retrieval Augmented Generation (RAG) (Lewis et al., 2020), see Figure 1.
+As noted by Wei et al.
+(2022a), training LLMs using a prompt packed with query-related information can yield substantial performance enhancement.
 Users can ask questions in the chat box by typing or using voice input.
 The response is displayed at the top of the page, and the document chunks used for the answer can be checked at the bottom (see Figure 2).
-2.0.1. Tool Dependencies
+3.1. Tool Dependencies
 The tool was constructed utilizing two innovative technologies - Gradio and LlamaIndex.
@@
 LlamaIndex, initially known as GPT Index, is a cutting-edge data framework designed for the efficient handling and accessibility of private or domain-specific data in LLMs applications.
-2.0.2. Knowledge Base Construction
+Since the factory documents can be long, they may overflow the LLM’s context window or result in unnecessary computational demand.
+To overcome this, we segment the materials into manageable chunks, each comprising approximately 400 tokens.
+This method effectively incorporates the materials into the LLM prompt without compromising the conversation flow.
+Following the segmentation, each document chunk is processed through LlamaIndex using the OpenAI Embedding API222https://api.openai.com/v1/embeddings—last accessed February 26, 2024.
+Utilizing the “text-embedding-ada-002” model, LlamaIndex transforms each chunk into a corresponding embedding vector.
+These resulting vectors are then securely stored, ready for future retrieval and use.
+3.2. Knowledge Base Construction
 Our experiment incorporates two distinct types of domain-specific data: factory manuals and shared knowledge from factory workers.
@@
 The materials come in various formats, including PDF, Word, and CSV files.
-To ensure easy access and indexing, we leverage the capabilities of LlamaIndex.
-In addition to the manuals, we integrate valuable practical knowledge from factory workers.
-This information is gathered from the production line, utilizing the five-why process, an iterative technique cited by (Serrat, 2017).
+In addition to the factory manuals, we integrate issue analysis reports from factory workers.
+This information is gathered from the production line, utilizing the five-why process, an iterative root-cause analysis technique (Serrat, 2017) (right side of Figure 2).
 The five-why technique probes into cause-and-effect relationships underlying specific problems by repeatedly asking ”Why?” until the root cause is revealed, typically by the fifth query.
 This process enables us to gather real-world issues encountered on production lines, which may not be covered in the factory manuals.
-Since these documents can be extensive, integrating them directly into the LLM conversation could be challenging.
-To overcome this, we further segment the materials into manageable chunks, each comprising approximately 400 tokens.
-This method effectively incorporates the materials into the LLM prompt without compromising the conversation flow.
-Following the segmentation, each document chunk is processed through LlamaIndex using the OpenAI Embedding API222https://api.openai.com/v1/embeddings.
-Utilizing the ‘text-embedding-ada-002’ model, LlamaIndex transforms each chunk into a corresponding embedding vector.
-These resulting vectors are then securely stored, ready for future retrieval and use.
-2.0.3. Query Construction
-To retrieve the document data relevant to specific user queries, we employ the same embedding model, “text-embedding-ada-002”, to generate vector representations of these queries.
+Upon entering all required information, including one or more “whys”, the operator presses “check”, triggering a prompt to the LLM that performs a logical check of the entered information and checks for inconsistencies with previously reported information.
+The operator can revise the entered information and submit it as is.
+Then, the submitted report will be added to a queue for expert operators to check before it is added to the knowledge base.
+3.3. Query Construction
+To retrieve the document data relevant to specific user queries, we employ the same embedding model, “text-embedding-ada-002” to generate vector representations of these queries.
 By leveraging the similarity calculation algorithm provided by LlamaIndex, we can identify and retrieve the top-K most similar segmented document snippets related to the user query.
 This allows us to construct pertinent LLM queries.
-Once the snippets are retrieved, they are synthesized into the following query template based on the templates used by LlamaIndex333https://docs.llamaindex.ai:
+Once the snippets are retrieved, they are synthesized into the following query template based on the templates used by LlamaIndex333https://docs.llamaindex.ai—last accessed February 26, 2024:
 You are an assistant that assists detergent production line operators with decision support and advice based on a knowledge base of standard operating procedures, single point lessons (SPL), etc.
@@
 This approach is designed to prevent potential user confusion from combining data from both sources into a single query.
-3. Model Benchmarking
-In our benchmarking experiment, we evaluated various commercial and open-source LLMs, including OpenAI’s ChatGPT (GPT-3.5 and GPT-4 from July 20th 2023), Guanaco 65B and 35B variants (Dettmers et al., 2023) based on Meta’s Llama (Large Language Model Meta AI) (Touvron et al., 2023a), Llama2 (Touvron et al., 2023b) with 70 billion parameters, and one of its derivatives, StableBeluga2 (Mahan et al., 2023).
-GPT-3, or Generative Pre-trained Transformer 3, is an autoregressive language model developed by OpenAI.
-This third-generation model utilizes deep learning to generate human-like text from a given prompt, capable of producing various types of data such as word lists or code lines (Floridi and Chiriatti, 2020).
-OpenAI recently employed a Reinforcement Learning from Human Feedback (RLHF) strategy to enhance GPT-3’s reliability, usefulness, and alignment (OpenAI, 2023).
-This technique relies on human labellers providing desired model behavior examples and ranking responses from the GPT-3 model to specific prompts (OpenAI, 2023).
-Using this feedback, OpenAI fine-tunes the GPT-3 model, creating ChatGPT, a leading language model, introduced as a prototype on November 30, 2022 (OpenAI, 2023).
-On the other hand, LLaMA, launched by Meta AI in February 2023, is a contrasting open-source language model (Touvron et al., 2023a).
-Meta AI trained a spectrum of models with parameters ranging from 7 billion to 65 billion.
-Like GPT-3, LLaMA also benefited from RLHF to enhance its capabilities.
-The LLaMA 7B model further paved the way for Alpaca’s development, a training regimen released by Taori et al.
-(2023).
-Alpaca employs a unique ”Self-Instruct” instruction tuning method, allowing it to achieve skills similar to OpenAI’s GPT-3 series (specifically the text-davinci-003 model) at a reduced cost.
-More recently, (Dettmers et al., 2023) released QLoRA, an efficient finetuning approach that allows 65B parameter models to be run on a single 48Gb GPU.
-These models, named Guanaco, are available in the following sizes: 7B, 13B, 33B, and 65B.
-According to the Vicuna benchmark, the 65B and 33B parameter models reached 99.3% and 97.3% ChatGPT performance, respectively.
-Nevertheless, like all LLaMA derivatives, they can only be used for research, not commercial applications.
-However, LLaMA 2 (Touvron et al., 2023b), which was released on July 18th, 2023, is available for both commercial and research use444https://about.fb.com/news/2023/07/llama-2/—last accessed January 10, 2024.
-We used a web UI for LLMs555https://github.com/oobabooga/text-generation-webui/tree/main—last accessed January 10, 2024 to load and test the Guanaco model and StableBeluga2 models.
-The models were loaded on a pair of Nvidia A6000s with NVlink and total Video Random Access Memory (VRAM) capacity of 96 GB.
+4. User Study in the Field
+We conducted a user study on the system to uncover perceived benefits, usability issues, risks, and barriers to adoption.
+The study comprised three tasks: to ask the system several questions as if they were operators, to fill in a “yellow tag” (issue analysis report) based on a recent issue and request a logical check, and finally, to upload new documents to the system.
+After each task, they were asked to provide feedback.
+Then, after completing all tasks, the participants were posed several open questions about the system’s benefits, risks, and barriers to adoption.
+Finally, demographic information, such as age, gender, and role, was collected.
+4.1. Participants
+We recruited participants from a detergent factory, of which were managers (P1-4), and were operators (P5-9).
+Of the nine participants, were women, and were men.
+Participant age was distributed over three brackets, namely were 30–39, were 40–49, and were 50–59.
+4.2. Qualitative Analysis
+An inductive thematic analysis (Guest et al., 2011) of the answers to the open questions resulted in six themes discussed below.
+Usability The theme of usability underlines the system’s ease of use and the need for clear instructions.
+Users mentioned the necessity for a “user-friendly” (P2) interface and highlighted the importance of having “more instructions and more details need to be loaded” (P1) to avoid confusion.
+This indicates a desire for intuitive navigation that could enable workers to use the system effectively without extensive training or referencing external help.
+The feedback suggests that the system already works well, as reflected in statements like “Easy-to-use system” (P3) and the system “works well” (P7).
+Access to information Users appreciated the “ease of having information at hand” (P1), facilitating immediate access to necessary documents.
+However, there is a clear call for improvements, such as the ability to “Include the possibility of opening IO, SPL, etc.
+in .pdf format for consultation” (P3).
+This theme is supported by requests for direct links to full documents, suggesting that while “the list of relevant documents from which information is taken is excellent” (P4), the ability to delve deeper into full documents would significantly enhance the user experience.
+Efficiency Users value the “greater speed in carrying out some small tasks” (P3).
+However, there are concerns about the system’s efficiency when it does not have the answer, leading to “wasting time looking for a solution to a problem in case it is not reported in the system’s history” (P3).
+Statements like “quick in responses” (P3) contrast with the need for questions to be “too specific to have a reliable answer” (P7), indicating tension between the desire for quick solutions and the system’s limitations.
+Adoption Users highlight several factors affecting adopting the new system.
+It includes challenges such as “awareness and training of operators [might hinder adoption]” (P3) and the need for “acceptance by all employees” (P4), which indicates that the system’s success is contingent on widespread user buy-in.
+The generational divide is also noted: “That older operators use it [on what may hinder adoption]” (P7) suggests that demographic factors may influence the acceptance of new technology.
+Safety Users express apprehension that “if the responses are not adequate, you risk safety” (P1), emphasizing the critical nature of reliable information in a factory setting.
+Moreover, the demand for updated and specific information underlines the importance of the system’s content being current and detailed to maintain operational safety standards, as stated by P9: “If it is updated and specific, it can help me”.
+Traditional versus Novel There is a noticeable preference for established practices among some users.
+For instance, “It’s faster and easier to ask an expert colleague working near me rather than [the system]” (P8) captures the reliance on human expertise over the assistant system.
+This tension is further demonstrated by the sentiment that “Operators may benefit more from traditional information retrieval systems” (P9), suggesting a level of skepticism or comfort with the status quo that the new system needs to overcome.
+5. LLM Benchmarking
+In our benchmarking experiment, we evaluated various commercial and open-source LLMs, including OpenAI’s ChatGPT (GPT-3.5 and GPT-4 from July 20th 2023), Guanaco 65B and 35B variants (Dettmers et al., 2023) based on Meta’s Llama (Large Language Model Meta AI) (Touvron et al., 2023a), Mixtral 8x7b (Jiang et al., 2024), Llama 2 (Touvron et al., 2023b), and one of its derivatives, StableBeluga2 (Mahan et al., 2023).
+This selection represents the state-of-the-art closed-sourced models (e.g., GPT-4) and open-source models (e.g., Llama 2).
+We included the (outdated) Guanaco models to demonstrate the improvements in the open-source sphere over the past year.
+We used a web UI for LLMs444https://github.com/oobabooga/text-generation-webui/tree/main—last accessed February 26, 2024 to load and test the Mixtral 8x7B, Guanaco models, and the StableBeluga2.
+The models were loaded on a pair of Nvidia A6000s with NVlink and a total Video Random Access Memory (VRAM) capacity of 96 GB.
 The 65B model was run in 8-bit mode to fit in the available VRAM.
-The models were used in the chat mode, with the llama-precise preset and fixed zero seed for reproducibility.
-Llama2 was evaluated using the demo on huggingface666https://huggingface.co/meta-llama/Llama-2-70b-chat-hf—last accessed January 10, 2024.
-To rigorously assess the models, we prepared 20 questions of varying complexity based on two types of context material: formal operating material and informal issue reports.
-The model prompt was constructed using the above template (2.0.3).
+We used the llama-precise parameter preset and fixed zero seed for reproducibility.
+Llama 2 was evaluated using the demo on huggingface555https://huggingface.co/meta-llama/Llama-2-70b-chat-hf—last accessed February 26, 2024.
+To rigorously assess the models, we prepared 20 questions of varying complexity based on two types of context material: half from operating manuals and half from unstructured issue reports.
+The operating manuals included excerpts from actual machine manuals and standard operating procedures, while the informal issue reports were free-text descriptions of issues we had previously collected from operators.
+The model prompt was constructed using the above template (3.3).
 Ultimately, the difficulty of a question is a combination of the question’s complexity and the clarity of the source material.
@@
 Conversely, difficult questions require more reasoning or comprise multiple parts, for example, ”What should I do if the central turntable is overloaded?” which has a nuanced answer dependent on several factors not clearly articulated in the context material.
-Every response is manually scored on factuality, completeness, and hallucinations.
-Factuality refers to aligning with the facts in the context material.
-Complete responses contain all the information relevant to the question in the context material.
-A hallucinated response appears grammatically and semantically coherent but is not based on the context material.
+In addition to measuring response length in words, every response is manually scored on factuality, completeness, and hallucinations as defined below:
+Factuality: Responses align with the facts in the context material.
+Completeness: Responses contain all the information relevant to the question in the context material.
+Hallucinations: Response appears grammatically and semantically coherent but is not based on the context material.
 The following scoring protocol is applied: one is awarded for a completely factual, complete, or hallucinated response.
@@
 The Guanaco models, based on Llama 1, perform significantly worse.
-The conciseness of the responses showed a similar pattern, except that StableBeluga2 produced the shortest answers (58 words), followed closely by GPT-4 (69 words).
+The conciseness of the responses showed a similar pattern, except that StableBeluga2 produced the shortest answers (58 words), followed closely by Mixtral 8x7B (66 words) and GPT-4 (69 words).
 | Model | Factuality | Completeness | Hallucinations | Words |
@@
 | StableBeluga2 | 95 | 92.5 | 7.5 | 58 |
+| Mixtral 8x7B | 92.5 | 92.5 | 2.5 | 66 |
 | GPT-3.5 | 90 | 90 | 5 | 89 |
@@
 | Guanaco 33b | 27.5 | 27.5 | 65.6 | 190 |
-4. User Study at the Factory
-We conducted a user study on the abovementioned system we co-designed with factory employees.
-We recruited factory managers for the user study.
-The study comprised three tasks: to ask the system several questions as if they were operators, to fill in a ”yellow tag” (issue report) based on a recent issue and request a logical check, and finally, to upload new documents to the system.
-After each task, they were asked to provide feedback.
-Then, they were posed with several open questions about benefits, risks, and challenges.
-Finally, demographic information was requested, such as age, gender, and role.
-An inductive thematic analysis of the answers to the open questions resulted in the following five themes:
-Usability (5): The system is praised for its user-friendliness, speed in delivering responses, and logical control function.
-Users appreciate the voice command and information search capabilities (P4).
-However, the user interface could be improved by having attachments appear after the main text for more accessible consultation (P1).
-Content (5): The system retrieves relevant information based on queries posed.
-Users appreciate access to relevant documents and seeing where the data is taken from.
-However, there’s a call for more detailed instructions to be added to the system.
-Also, it’s mentioned that the part of answers based on yellow tags sometimes refers to previous knowledge, causing confusion (P2).
-Features (4): There are suggestions to include additional features, such as the ability to open IO, SPL, and yellow tags in .pdf format for consultation (P3).
-Furthermore, one participant appreciated the ability to change file names (P4).
-Lastly, another participant could not access the yellow tag insertion fields as the fields were in a collapsed state (P1).
-Risks & Benefits (4): The system is recognized for its potential to modernize the factory and make specific tasks faster, for example, ”Greater speed in carrying out some small activities.” (P3) and ”Being able to easily find information, especially if they are in different documents.” (P4).
-However, concerns are raised about the safety risks if the system provides inadequate answers and the potential time-wasting if the system lacks information on specific problems (P2, P3).
-Employee Acceptance & Training (2): Users hope for system acceptance among all employees.
-There’s also a suggestion about training and sensitization of the operators to the system’s advantages (P3).
-P2 implies that users may turn to a human expert if available instead of using the system as the ”[Information from the system] might be less quick, less effective, and less exhaustive than asking the expert staff present.”
-5. Summary and Discussion
-5.1. Model Benchmarking
-GPT-4 performs best across all measures but is closely followed by StableBeluga2 and GPT-3.5.
-Compared to GPT-4, the cost per input token for GPT-3.5 is significantly lower777https://openai.com/pricing#language-models—last accessed January 10, 2024.
-The higher costs of GPT-4 are partially counteracted by the response length as GPT-3.5 (and Llama 2) tended to be wordier and include additional details that were not directly requested, whereas GPT-4 (and StableBeluga2) generated more concise responses.
-The latest generation of open-source models based on Llama-2 demonstrates a clear jump forward relative to its predecessors based on Llama-1, which were more prone to hallucinations and exhibited poorer reasoning abilities over the context material.
-While open-source models like StableBeluga2 do not score as high as GPT-4, they ensure better data security, privacy, and customization.
-Unlike GPT-4 and GPT-3.5, they can be hosted locally and heavily customized.
+6. Discussion
+6.1. GPT-4 is the best, but open-source models follow closely
+GPT-4 performs best across all measures but is closely followed by StableBeluga2, Mixtral 8x7B, and GPT-3.5.
+Compared to GPT-4, the cost per input token for GPT-3.5 is significantly lower666https://openai.com/pricing#language-models—last accessed February 26, 2024.
+However, the higher costs of GPT-4 are partially counteracted by its concise yet complete responses.
+If longer, more detailed responses were desired (e.g., for training purposes), the prompt could be adjusted.
+We observed that the less powerful models, such as GPT-3.5 and Llama 2, tended to be wordier and include additional details that were not directly requested.
+In contrast, GPT-4, StableBeluga2, and Mixtral 8x7B generated more concise responses.
+The latest generation of open-source models, such as Mixtral 8x7B and Llama 2 variants, such as StableBeluga2, demonstrates a clear jump forward relative to their predecessors based on Llama-1, which were more prone to hallucinations and exhibited poorer reasoning abilities over the context material.
+While open-source models like StableBeluga2 and Mixtral 8x7B do not score as high as GPT-4, they ensure better data security, privacy, and customization if hosted locally.
 This can be a crucial consideration for companies with sensitive data or unique needs.
-It should be noted that fine-tuning GPT-3.5 is possible from August 22nd 2023, and it was announced that fine-tuning GPT-4 would be available in the fall of 2023 888https://openai.com/blog/gpt-3-5-turbo-fine-tuning-and-api-updates—last accessed January 10, 2024.
-5.2. Perceptions at the Factory
+6.2. The tool is beneficial but inferior to human experts
 Users appreciate the system’s functionality and see it as a tool for modernizing factory operations and speeding up operations.
@@
 However, they express concerns about potential safety risks and the efficacy of information retrieval compared to consulting expert personnel.
-Although the tool is seen as inferior to learning from a human expert, it can help when human experts are not present or are unfamiliar with the issue at hand.
-5.3. Limitations
-The participant pool for the user study comprised of factory managers and was limited in size.
-This potentially overlooks the varied perspectives of other stakeholders, such as factory operators.
-To partially mitigate this we asked the participants to pose queries as if they were operators.
+While these concerns are understandable, the tool was not designed to replace human-human interactions; instead, it can be used when no human experts are present or when they do not know or remember how to solve a specific issue.
+This would come into play during the night shift at the factory where we conducted the user study as a single operator operates a production line, leaving limited options for eliciting help from others.
+6.3. Limitations and future work
 We used the same prompt for all LLMs; however, it is possible that some of the LLMs would perform better with a prompt template developed explicitly for it.
-Furthermore, we matched the settings of the LLMs (e.g., temperature) as closely as possible across all the tested models; however, the same settings across model types were not completely equivalent, and in the case of Llama 2, we did not have access to the presets as we did not host it locally.
-The study’s design did not include a comprehensive real-world evaluation involving end users in their natural working environment, as this was considered too risky for our industry partner.
-Such an environment might present unique challenges and considerations not addressed in this research.
-Our benchmarking procedure involved only 20 questions, and a singular coder assessed the responses.
-This introduces the potential for bias and limits the generalizability of our findings.
-Furthermore, the limited number of questions may not cover the full spectrum of complexities in real-world scenarios.
-These limitations suggest directions for future research, where a broader participant pool, more comprehensive prompt and model customization, and real-world evaluations could be considered.
-Furthermore, considering the pace at which LLMs are improving, it would be desirable to (semi)automate benchmarking.
+However, we matched the LLMs’ hyperparameters (e.g., temperature) as closely as possible across all the tested models, except for Llama 2, as we did not have access to the presets as we did not host it locally.
+Our model benchmarking procedure involved 20 questions, and a singular coder assessed the responses.
+This introduces the potential for bias, and the limited number of questions may not cover the full spectrum of complexities in real-world scenarios.
+However, we varied query complexity and source material types to (partially) mitigate these shortcomings.
+The study’s design did not include a real-world evaluation involving end users operating the production line, as this was considered too risky for our industry partner.
+Such an environment might present unique challenges and considerations not addressed in this research, such as time pressure.
+However, by involving operators and managers and instructing them to pose several questions based on their actual work experience, we could still evaluate the system and collect valid feedback.
+These limitations suggest directions for future research, for example, longitudinal studies where operators use the tool during production line operations and more comprehensive prompt and model customization.
+Longitudinal studies will be key to understanding the real-world impact on production performance, operator well-being, and cognitive abilities.
 7. Conclusion
 The results demonstrated GPT-4’s superior performance over other models regarding factuality, completeness, and minimal hallucinations.
-Interestingly, open source models like StableBeluga2 followed close behind.
+Interestingly, open-source models like StableBeluga2 and Mixtral 8x7B followed close behind.
 The user study highlighted the system’s user-friendliness, speed, and logical functionality.
 However, improvements in the user interface and content specificity were suggested, along with potential new features.
-Benefits included modernizing factory operations and speeding up certain tasks, though concerns about safety and efficiency were raised.
+Benefits included modernizing factory operations and speeding up specific tasks, though concerns about safety, efficiency, and inferiority to asking human experts were raised.
 Acknowledgements.
@@
 - (1)
+- Man (2014) 2014.
+The Manufacturing Knowledge Repository - Consolidating Knowledge to Enable Holistic Process Knowledge Management in Manufacturing:.
+In Proceedings of the 16th International Conference on Enterprise Information Systems (Lisbon, Portugal).
+SCITEPRESS - Science and and Technology Publications, 39–51.
+https://doi.org/10.5220/0004891200390051
 - Abid et al.
@@
 Volume 87-Data Mining and Analytics 2008 (2008), 193–199.
-- Floridi and Chiriatti (2020) Luciano Floridi and Massimo Chiriatti. 2020. GPT-3: Its nature, scope, limits, and consequences. Minds and Machines 30, 4 (2020), 681–694.
+- Fantini et al.
+(2020) Paola Fantini, Marta Pinzone, and Marco Taisch.
+2020.
+Placing the operator at the centre of Industry 4.0 design: Modelling and assessing human activities within cyber-physical systems.
+Computers & Industrial Engineering 139 (2020), 105058.
+https://doi.org/10.1016/j.cie.2018.01.025
+- Feng et al.
+(2017) Shaw C.
+Feng, William Z.
+Bernstein, Jr.
+Thomas Hedberg, and Allison Barnard Feeney.
+2017.
+Toward Knowledge Management for Smart Manufacturing.
+Journal of computing and information science in engineering 17 3 (2017).
+https://api.semanticscholar.org/CorpusID:4780425
 - Gao et al.
@@
 https://doi.org/10.18653/v1/2021.acl-long.295
+- Guest et al. (2011) Greg Guest, Kathleen M MacQueen, and Emily E Namey. 2011. Applied thematic analysis. sage publications.
 - Jang et al.
@@
 arXiv:2212.14882 [cs.CL]
+- Jiang et al.
+(2024) Albert Q.
+Jiang, Alexandre Sablayrolles, Antoine Roux, Arthur Mensch, Blanche Savary, Chris Bamford, Devendra Singh Chaplot, Diego de las Casas, Emma Bou Hanna, Florian Bressand, Gianna Lengyel, Guillaume Bour, Guillaume Lample, Lélio Renard Lavaud, Lucile Saulnier, Marie-Anne Lachaux, Pierre Stock, Sandeep Subramanian, Sophia Yang, Szymon Antoniak, Teven Le Scao, Théophile Gervet, Thibaut Lavril, Thomas Wang, Timothée Lacroix, and William El Sayed.
+2024.
+Mixtral of Experts.
+arXiv:2401.04088 [cs.LG]
 - Kernan Freire et al.
-(2023) Samuel Kernan Freire, Mina Foosherian, Chaofan Wang, and Evangelos Niforatos.
-2023.
+(2023a) Samuel Kernan Freire, Mina Foosherian, Chaofan Wang, and Evangelos Niforatos.
+2023a.
 Harnessing Large Language Models for Cognitive Assistants in Factories.
@@
 https://doi.org/10.1145/3571884.3604313
+- Kernan Freire et al.
+(2023b) Samuel Kernan Freire, Chaofan Wang, Santiago Ruiz-Arenas, and Evangelos Niforatos.
+2023b.
+Tacit Knowledge Elicitation for Shop-floor Workers with an Intelligent Assistant.
+In Extended Abstracts of the 2023 CHI Conference on Human Factors in Computing Systems.
+1–7.
+https://doi.org/10.1145/3544549.3585755
 - Kwon and Mihindukulasooriya (2022) Bum Chul Kwon and Nandana Mihindukulasooriya.
@@
 https://doi.org/10.18653/V1/2022.TRUSTNLP-1.7
-- Lahat et al.
-(2023) Adi Lahat, Eyal Shachar, Benjamin Avidan, Zina Shatz, Benjamin S.
-Glicksberg, and Eyal Klang.
-2023.
-Evaluating the use of large language model in identifying top research questions in gastroenterology.
-Scientific Reports 2023 13:1 13 (3 2023), 1–6.
-Issue 1.
-https://doi.org/10.1038/s41598-023-31412-2
+- Leoni et al.
+(2022) Luna Leoni, Marco Ardolino, Jamal El Baz, Ginetta Gueli, and Andrea Bacchetti.
+2022.
+The Mediating Role of Knowledge Management Processes in the Effective Use of Artificial Intelligence in Manufacturing Firms.
+42, 13 (2022), 411–437.
+https://doi.org/10.1108/IJOPM-05-2022-0282
 - Lewis et al.
@@
 - Maddikunta et al.
-(2022) Praveen Kumar Reddy Maddikunta, Quoc-Viet Pham, Prabadevi B, N Deepa, Kapal Dev, Thippa Reddy Gadekallu, Rukhsana Ruby, and Madhusanka Liyanage.
-2022.
+(2022a) Praveen Kumar Reddy Maddikunta, Quoc-Viet Pham, Prabadevi B, N Deepa, Kapal Dev, Thippa Reddy Gadekallu, Rukhsana Ruby, and Madhusanka Liyanage.
+2022a.
 Industry 5.0: A survey on enabling technologies and potential applications.
@@
 https://doi.org/10.1016/j.jii.2021.100257
+- Maddikunta et al.
+(2022b) Praveen Kumar Reddy Maddikunta, Quoc-Viet Pham, B Prabadevi, Natarajan Deepa, Kapal Dev, Thippa Reddy Gadekallu, Rukhsana Ruby, and Madhusanka Liyanage.
+2022b.
+Industry 5.0: A survey on enabling technologies and potential applications.
+Journal of Industrial Information Integration 26 (2022), 100257.
 - Mahan et al.
@@
 [https://huggingface.co/stabilityai/StableBeluga2](https://huggingface.co/stabilityai/StableBeluga2)
+- May et al.
+(2015) Gökan May, Marco Taisch, Andrea Bettoni, Omid Maghazei, Annarita Matarazzo, and Bojan Stahl.
+2015.
+A New Human-centric Factory Model.
+26 (2015), 103–108.
+https://doi.org/10.1016/j.procir.2014.07.112
 - Mercedes-Benz (2023) Mercedes-Benz.
@@
 - Nov et al. (2023) Oded Nov, Nina Singh, and Devin Mann. 2023. Putting ChatGPT’s Medical Advice to the (Turing) Test. arXiv:2301.10035 [cs.HC]
-- OpenAI (2023) OpenAI. 2023. ChatGPT: Optimizing Language Models for Dialogue. https://openai.com/blog/chatgpt/.
+- Richter et al.
+(2019) Shahper Richter, Lena Waizenegger, Melanie Steinhueser, and Alexander Richter.
+2019.
+Knowledge Management in the Dark: The Role of Shadow IT in Practices in Manufacturing.
+15, 2 (2019), 1–19.
+https://doi.org/10.4018/IJKM.2019040101
 - Semnani et al. (2023) Sina J. Semnani, Violet Z. Yao, Heidi C. Zhang, and Monica S. Lam. 2023. WikiChat: A Few-Shot LLM-Based Chatbot Grounded with Wikipedia. arXiv:2305.14292 [cs.CL]
 - Serrat (2017) Olivier Serrat. 2017. The five whys technique. Knowledge solutions: Tools, methods, and approaches to drive organizational performance (2017), 307–310.
+- Shneiderman (2022) Ben Shneiderman. 2022. Human-centered AI. Oxford University Press.
 - Singhal et al.
@@
 - Tang et al. (2023) Ruixiang Tang, Xiaotian Han, Xiaoqian Jiang, and Xia Hu. 2023. Does Synthetic Data Generation of LLMs Help Clinical Text Mining? arXiv:2303.04360 [cs.CL]
-- Taori et al.
-(2023) Rohan Taori, Ishaan Gulrajani, Tianyi Zhang, Yann Dubois, Xuechen Li, Carlos Guestrin, Percy Liang, and Tatsunori B.
-Hashimoto.
-2023.
-Stanford Alpaca: An Instruction-following LLaMA model.
-https://github.com/tatsu-lab/stanford_alpaca.
 - Touvron et al.
@@
 - Wang et al.
-(2023) Xingzhi Wang, Nabil Anwer, Yun Dai, and Ang Liu.
-2023.
+(2023a) Xingzhi Wang, Nabil Anwer, Yun Dai, and Ang Liu.
+2023a.
 ChatGPT for design, manufacturing, and education.
@@
 https://doi.org/10.1016/j.procir.2023.04.001 The 33rd CIRP Design Conference.
+- Wang et al.
+(2023b) Zezhong Wang, Fan Yang, Pu Zhao, Lu Wang, Jue Zhang, Mohit Garg, Qingwei Lin, and Dongmei Zhang.
+2023b.
+Empower Large Language Model to Perform Better on Industrial Domain-Specific Question Answering.
+ArXiv abs/2305.11541 (2023).
+https://api.semanticscholar.org/CorpusID:258822888
 - Wei et al. (2023) Chengwei Wei, Yun-Cheng Wang, Bin Wang, and C. C. Jay Kuo. 2023. An Overview on Language Models: Recent Developments and Outlook. (3 2023). https://arxiv.org/abs/2303.05759v1
 - Wei et al.
-(2022) Jason Wei, Yi Tay, Rishi Bommasani, Colin Raffel, Barret Zoph, Sebastian Borgeaud, Dani Yogatama, Maarten Bosma, Denny Zhou, Donald Metzler, Ed H.
+(2022a) Jason Wei, Yi Tay, Rishi Bommasani, Colin Raffel, Barret Zoph, Sebastian Borgeaud, Dani Yogatama, Maarten Bosma, Denny Zhou, Donald Metzler, Ed H.
 Chi, Tatsunori Hashimoto, Oriol Vinyals, Percy Liang, Jeff Dean, and William Fedus.
-2022.
+2022a.
 Emergent Abilities of Large Language Models.
 arXiv:2206.07682 [cs.CL]
+- Wei et al.
+(2022b) Jason Wei, Xuezhi Wang, Dale Schuurmans, Maarten Bosma, brian ichter, Fei Xia, Ed Chi, Quoc V Le, and Denny Zhou.
+2022b.
+Chain-of-Thought Prompting Elicits Reasoning in Large Language Models.
+In Advances in Neural Information Processing Systems, S.
+Koyejo, S.
+Mohamed, A.
+Agarwal, D.
+Belgrave, K.
+Cho, and A.
+Oh (Eds.), Vol.
+35.
+Curran Associates, Inc., 24824–24837.
+https://proceedings.neurips.cc/paper_files/paper/2022/file/9d5609613524ecf4f15af0f7b31abca4-Paper-Conference.pdf
 - Wellsandt et al.
@@
 Springer International Publishing, Cham, 321–330.
+- Wen et al.
+(2023) Cheng Wen, Xianghui Sun, Shuaijiang Zhao, Xiaoquan Fang, Liang Chen, and Wei Zou.
+2023.
+ChatHome: Development and Evaluation of a Domain-Specific Language Model for Home Renovation.
+ArXiv abs/2307.15290 (2023).
+https://api.semanticscholar.org/CorpusID:260315830
 - Xia et al.
@@
 - Xie et al.
-(2023) Qianqian Xie, Weiguang Han, Xiao Zhang, Yanzhao Lai, Min Peng, Alejandro Lopez-Lira, and Jimin Huang.
-2023.
+(2023a) Qianqian Xie, Weiguang Han, Xiao Zhang, Yanzhao Lai, Min Peng, Alejandro Lopez-Lira, and Jimin Huang.
+2023a.
 PIXIU: A Large Language Model, Instruction Data and Evaluation Benchmark for Finance.
@@
 https://arxiv.org/abs/2306.05443v1
+- Xie et al.
+(2023b) Tong Xie, Yuwei Wan, Wei Huang, Zhenyu Yin, Yixuan Liu, Shaozhou Wang, Qingyuan Linghu, Chunyu Kit, Clara Grazian, Wenjie Zhang, Imran Razzak, and Bram Hoex.
+2023b.
+DARWIN Series: Domain Specific Large Language Models for Natural Science.
+ArXiv abs/2308.13565 (2023).
+https://api.semanticscholar.org/CorpusID:264960172
 - Xu et al.
@@
 - Xu et al.
-(2021) Xun Xu, Yuqian Lu, Birgit Vogel-Heuser, and Lihui Wang.
-2021.
+(2021a) Xun Xu, Yuqian Lu, Birgit Vogel-Heuser, and Lihui Wang.
+2021a.
+Industry 4.0 and Industry 5.0—Inception, conception and perception.
+Journal of Manufacturing Systems 61 (2021), 530–535.
+- Xu et al.
+(2021b) Xun Xu, Yuqian Lu, Birgit Vogel-Heuser, and Lihui Wang.
+2021b.
 Industry 4.0 and Industry 5.0—Inception, conception and perception.
@@
 - Zhang et al.
-(2023) Jianzhang Zhang, Yiyang Chen, Nan Niu, and Chuang Liu.
-2023.
+(2023a) Jianzhang Zhang, Yiyang Chen, Nan Niu, and Chuang Liu.
+2023a.
 A Preliminary Evaluation of ChatGPT in Requirements Information Retrieval.
@@
 https://arxiv.org/abs/2304.12562v1
+- Zhang et al.
+(2023b) Wenxuan Zhang, Hongzhi Liu, Yingpeng Du, Chen Zhu, Yang Song, Hengshu Zhu, and Zhonghai Wu.
+2023b.
+Bridging the Information Gap Between Domain-Specific Model and General LLM for Personalized Recommendation.
+ArXiv abs/2311.03778 (2023).
+https://api.semanticscholar.org/CorpusID:265043714
 - Zhao et al.
@@
 arXiv:2303.18223 [cs.CL]
+- Zuccon et al.
+(2023) Guido Zuccon, Bevan Koopman, and Razia Shaik.
+2023.
+ChatGPT Hallucinates when Attributing Answers.
+In Proceedings of the Annual International ACM SIGIR Conference on Research and Development in Information Retrieval in the Asia Pacific Region (¡conf-loc¿, ¡city¿Beijing¡/city¿, ¡country¿China¡/country¿, ¡/conf-loc¿) (SIGIR-AP ’23).
+Association for Computing Machinery, New York, NY, USA, 46–51.
+https://doi.org/10.1145/3624918.3625329